OpenStudy (anonymous):

5. A bag contains 6 silver and 5 gold coins. Which probability distribution represents the number of silver coins in a random draw of 4 coins?

OpenStudy (anonymous):

ans

OpenStudy (kropot72):

This is sampling without replacement which is represented by the hypergeometric distribution.

OpenStudy (anonymous):

explain i am not getting

OpenStudy (kropot72):

The hypergeometric distribution is generated when a sample of size n is taken, without replacement from a population containing a items of one type and b items of others. \[p(x)=\frac{\left(\begin{matrix}a \\ x\end{matrix}\right)\left(\begin{matrix}b \\ n-x\end{matrix}\right)}{\left(\begin{matrix}a+b \\ n\end{matrix}\right)}\]